A rectangular lawn measures 8 m by 4 m and is surrounded by a flowerbed of
uniform width. The combined area of the lawn and flowerbed is 6 more than
triple the area of the lawn alone. What is the width of the flowerbed?

Answer by Boreal(15235)

Lawn with area of 32m^2
Combined area is 96 m^2.
the length of the combined area is 8 m +2 x m
the width is 4 m+2 x m
the area of the combined is 32+24x+4x^2=102
so 4x^2+24x-70=0
2x^2+12x-35=0
x=(1/4)(-12 +/- sqrt (144+280)); sqrt 424=20.59
positive root is 2.15 m the width
-
dimensions are 12.3m x 8.3m=102.09 m (rounding error)